# Here Be Dragons: Bimodal posteriors arise from numerical integration error in longitudinal models

### v1

@@ 1.2 What’s going on inside the numerical method?
The affine DE with negative slope always converges to the asymptotic size $\hat{\beta}_{0}/\hat{\beta}_{1}=\alpha$, and in fact if we make the substitution $\alpha\hat{\beta}_{1}=\hat{\beta}_{0}$ we can refine further to get
$0=Y(t_{j})-Y(t_{j}+h)+(\alpha-Y(t_{j}))\bigg{(}h\hat{\beta}_{1}-\frac{h^{2}}{2}\hat{\beta}_{1}^{2}+\frac{h^{3}}{6}\hat{\beta}_{1}^{3}-\frac{h^{4}}{24}\hat{\beta}_{1}^{4}\bigg{)}$ (5)
in which constants are $Y(t_{j})$, $Y(t_{j}+h)$, $t_{j}$, $h$, and $\alpha$. Given these values, the $\hat{\beta}_{1}$ solutions of Equation (5) will be the non-identifiable posterior modes. Through further work it may be possible to prove general non-identifiability for the von Bertalanffy model (affine first order ODE with negative slope) under RK4 with fixed step size. Such a proof is outside the scope of this thesis.
In our example case the asymptotic solution is $\hat{\beta}_{0}/\hat{\beta}_{1}=10=\alpha$ (this is a very strong relationship in the data too), so we get
$0=Y(t_{j})-Y(t_{j}+h)+(10-Y(t_{j}))\bigg{(}h\hat{\beta}_{1}-\frac{h^{2}}{2}\hat{\beta}_{1}^{2}+\frac{h^{3}}{6}\hat{\beta}_{1}^{3}-\frac{h^{4}}{24}\hat{\beta}_{1}^{4}\bigg{)}.$
In theory Equation (5) could be expanded to include all the time-steps, iterating through substitutions of $Y(t_{j}+2h)$ and so on. Any values of $\hat{\beta}_{1}$ that solved the extended version of Equation (5) would be indistinguishable in the posterior. Such a calculation is outside the scope of this paper, but we can look at the solutions for each time step. We implemented Equation (5) in R and looked at the first five steps of $h=0.5$ from Figure 1 to see if there are multiple zeroes. The results are in Figure 2 and show that each step has two solutions, and that those solutions are the same across the steps.
We can also do an explicit calculation for the first step with our estimated parameters. For our true values $\beta_{0}=10$, $\beta_{1}=1$ and $Y(t_{j})=Y(0)=1$, the analytic solution of the first step given $h=1/2$ is $Y(0.5)=4.541224$. Substitution of $h$, $Y(t_{j})$, and $Y(t_{j}+h)$ into Equation (5) gives us a formula for the first step:
$0=1-4.541224+9\bigg{(}\frac{1}{2}\hat{\beta}_{1}-\frac{1}{8}\hat{\beta}_{1}^{2}+\frac{1}{48}\hat{\beta}_{1}^{3}-\frac{1}{384}\hat{\beta}_{1}^{4}\bigg{)}.$ (6)
The two solutions to Equation (6) are $\beta_{1}=1.0008$ and $4.9112$, which aligns with Figure 2 and what we observe in our posterior bimodality (allowing for numerical error and noise).
To show that this problem is not a quirk of our chosen parameter values, we consider Equation (5) for $\beta_{0}=8$, $\beta_{1}=1.5$, $Y(t_{j})=1$. The asymptotic size is $\beta_{0}/\beta_{1}=\alpha=16/3$, so we substitute in $\hat{\beta}_{0}=\frac{16}{3}\hat{\beta}_{1}$ and get
$0=Y(t_{j})-Y(t_{j}+h)+\frac{13}{3}\bigg{(}\frac{1}{2}\hat{\beta}_{1}-\frac{1}{8}\hat{\beta}_{1}^{2}+\frac{1}{48}\hat{\beta}_{1}^{3}-\frac{1}{384}\hat{\beta}_{1}^{4}\bigg{)}.$ (7)
The solutions to Equation (7) for $Y(t_{j})=1$ and $h=0.5$ are $\hat{\beta}_{1}=1.5076$ and $4.5678$, and we have non-identifiability again.
@@
Equation 5 is calculable directly because we have an analytic solution, explicit numerical method, and a straight-forward function. Attempts to do the same with the non-linear Canham function (Canham et al.,, 2004) for example would be impossible as there is no analytic solution for comparison. While the substitution with the analytic solution is not essential to testing a step where a better numerical method can give $Y(t_{j}+h)$, it would be intractable for Canham anyway due to the non-linear DE. In general it may not be possible to do an explicit formula of that form. Simulation is much easier and the user can test different methods and estimation tools as well.
2 Methods
We have three questions about the pathological behaviour, each of which required different testing.
@@ 2 Methods
  • Do our chosen solutions give unimodal, unbiased estimates? We do simulation studies for the affine and Canham models to test whether the solutions implemented in hmde have reliable results.
The data we use is simulated so we know the true values. We produce ‘true’ data from the analytic solution Equation (3) with parameters $\beta_{0}=10$, $\beta_{1}=1$ and initial size $Y(0)=1$. The simulated survey structure is 10 size values $Y(t_{j})$ with an interval of 1 unit of time between them, so $t_{j+1}-t_{j}=1$. From the analytic sizes over time we get ‘observations’ with error, rounded to precision 0.1, of the form
@@
For each test we ran 10,000 fits. In the case of MCMC this is 10,000 independent chains, while for the optimisation algoithm it is 10,000 runs each of which randomly selects a starting point in the parameter space. The large sample size was chosen because in initial testing we found that the second mode for the smallest RK4 step size was very rare and wanted to effectively estimate the mixture model probability.
@@
From each chain we extract the estimates $\hat{\beta}_{0}$ and $\hat{\beta}_{1}$ as the means of the posterior samples. We fit multivariate normal finite mixture models (McLachlan et al.,, 2019) to the 10,000 estimates for each MCMC model using the mixtools package (Benaglia et al.,, 2009). We can identify the posteriors that have two clusters through visual analysis of scatter plots as the clusters are very distant from each other, and the distance is much larger than the within-cluster variance with no values in the space between them. To initialise the finite mixture algorithm, we classify clusters by taking points on either side of the sample mean for $\hat{\beta}_{0}$, giving
@@
We initialise the finite mixture algorithm with the sample means of the two groups to avoid convergence problems. The fitted mixture of normal distributions is then used to analyse the clustering behaviour, as we get cluster mean estimate vectors, covariance matrices, and cluster occurrence probability estimates. We also use scatter plots to look at the cluster marginal posterior parameter distributions. In the unimodal case we look at the mean of the estimated posterior parameters to understand bias, and a 95% posterior credible interval (Gelman et al.,, 2021) built from taking the central 95% quantile interval of the posterior estimates to get an idea of spread.
With the estimates in hand we can independently verify the numerical error using the ode function from the deSolve package (Karline Soetaert et al.,, 2010), which has its own implementation of the Runge-Kutta 4th order method (rk4). We plot the $Y(t_{j})$ values for the true and estimated parameters with the step size used in the MCMC sampling and show that they are statistically indistinguishable by calculating the root mean squared error. We use plots of the analytic solutions to show that the erroneous parameter combinations actually produce very different behaviour in the absence of numerical error.

### v2

@@ 2.1 Numerical integration theory
The affine DE with negative slope always converges to the asymptotic size $\beta_{0}/\beta_{1}=\alpha$, and in fact if we make the substitution $\alpha\beta_{1}=\beta_{0}$ we can reduce to a function of a single variable $\beta_{1}$ to get
$\hat{Y}(t_{j}+h)=Y(t_{j})+(\alpha-Y(t_{j}))\bigg{(}h\beta_{1}-\frac{h^{2}}{2}\beta_{1}^{2}+\frac{h^{3}}{6}\beta_{1}^{3}-\frac{h^{4}}{24}\beta_{1}^{4}\bigg{)}$
as required. ∎
In the longitudinal model, we are estimating $\beta_{1}$ values that minimise the difference between sizes over time based on the numerical projection using estimated parameters and the data arising from the analytic solution. In a zero error situation, this involves finding $\hat{\beta}_{1}$ values which produce numerical steps that match the analytic solution. That is, minimising the difference between $Y(t_{j}+h)-Y(t_{j})$ and $\hat{Y}(t_{j}+h)-Y(t_{j})$, where the former uses the true parameter $\beta_{1}$ and the analytic solution to reach $Y(t_{j}+h)$ while the latter uses the estimated parameter $\hat{\beta}_{1}$ and the numerical approximation. We need to distinguish the true parameter $\beta_{1}$ as used in the analytic solution from what we are estimating – $\hat{\beta}_{1}$ – which only forms part of the numerical step, hence the hat. We are assuming that both start at the same true $Y(t_{j})$. Lemma 2.1 gives us an algebraic form for one step, but we want a general form for the full longitudinal model with $n$ steps.
Proposition 2.2.
For an affine first order ODE of the form
$\frac{dY}{dt}=\beta_{0}-\beta_{1}Y(t)$
within a longitudinal model using $n$ steps of the Runge-Kutta 4th order algorithm, constrain estimates to fall along the line $\hat{\beta}_{0}=\alpha\hat{\beta}_{1}$. Then $\hat{\beta}_{1}$ solves
$10=h\hat{\beta}_{1}-\frac{h^{2}}{2}\hat{\beta}_{1}^{2}+\frac{h^{3}}{6}\hat{\beta}_{1}^{3}-\frac{h^{4}}{24}\hat{\beta}_{1}^{4}+e^{-\beta_{1}h}-$ 1 . (2.8)
Proof.
We start by looking at what happens for a single step, so set $n=1$. From Lemma 2.1 we have
$\hat{Y}(t_{j}+h)=Y(t_{j})+(\alpha-Y(t_{j}))f(\hat{\beta}_{1},h)$
where
$f(\hat{\beta}_{1},h)=h\hat{\beta}_{1}-\frac{h^{2}}{2}\hat{\beta}_{1}^{2}+\frac{h^{3}}{6}\hat{\beta}_{1}^{3}-\frac{h^{4}}{24}\hat{\beta}_{1}^{4}$
for convenience, as it will not change throughout this proof. Values of $\hat{\beta}_{1}$ estimated on data will be those that match the numerical solution to the analytic one, giving
$Y(t_{j}+h)-Y(t_{j})=(\alpha-Y(t_{j}))f(\hat{\beta}_{1},h)$ (2.9)
where the left hand size is the true growth increment between $t_{j}$ and $t_{j}+h$ arising from the analytic solution which uses the true parameter $\beta_{1}$, and the right is the numerical estimate of the increment that depends on the estimate $\hat{\beta}_{1}$.
From the analytic solution in Equation (2.3) we know that
$Y(t_{j}+h)=\alpha+(Y(t_{j})-\alpha)e^{-\beta_{1}(t_{j}+h-t_{j})}=\alpha+(Y(t_{j})-\alpha)e^{-\beta_{1}h},$
which we can substitute into Equation (2.9) to get
$(\alpha-Y(t_{j}))f(\hat{\beta}_{1},h)=$ $\ \alpha+(Y(t_{j})-\alpha)e^{-\beta_{1}h}-Y(t_{j})$
$=$ $\ (Y(t_{j})-\alpha)e^{-\beta_{1}h}-(Y(t_{j})-\alpha)$
$=$ $\ (Y(t_{j})-\alpha)(e^{-\beta_{1}h}-1).$ (2.10)
When solving for $\hat{\beta}_{1}$ we re-arrange Equation (2.10) to get
$0=$ $\ (\alpha-Y(t_{j}))f(\hat{\beta}_{1},h)-(Y(t_{j})-\alpha)(e^{-\beta_{1}h}-1)$
$=$ $\ (\alpha-Y(t_{j}))\big{(}f(\hat{\beta}_{1},h)+e^{-\beta_{1}h}-1\big{)}.$ (2.11)
The only variable in Equation (2.11) is $\hat{\beta}_{1}$, so the solutions will be the same as those that solve
$f(\hat{\beta}_{1},h)+e^{-\beta_{1}h}-1=0.$ (2.12)
The $n$-step case comes into play when we have a sequence of $Y(t_{j})$ that we are fitting to, or are dividing up the time between two and applying RK4 repeatedly. Repeated applications of the algorithm is one way of seeking greater numerical accuracy through a smaller step size. Taking from Equation (2.1), but extending to $n$ steps gives
$\hat{Y}(t_{j}+nh)=Y(t_{j})$ $+\underbrace{(\alpha-Y(t_{j}))f(\hat{\beta}_{1},h)}_{\text{step 1}}$
$+\underbrace{(\alpha-Y(t_{j}+h))f(\hat{\beta}_{1},h)}_{\text{step2}}$
$+\ldots+\underbrace{(\alpha-Y(t_{j}+(n-1)h))f(\hat{\beta}_{1},h)}_{\text{step n}}$
$=Y(t_{j})$ $+f(\hat{\beta}_{1},h)\Bigg{(}\sum_{i=0}^{n-1}\alpha-Y(t_{j}+ih)\Bigg{)}.$
As in the 1-step case, the $n$-step model will seek $\hat{\beta}_{1}$ values such that $\hat{Y}(t_{j}+nh)=Y(t_{j}+nh)$ to match the data, so we want
$Y(t_{j}+nh)-Y(t_{j})=f(\hat{\beta}_{1},h)\Bigg{(}\sum_{i=0}^{n-1}\alpha-Y(t_{j}+ih)\Bigg{)}.$ (2.13)
As in the 1-step case we take the left hand side (LHS) of Equation (2.13) and substitute in the analytic solution $Y(t_{j}+nh)=\alpha+(Y(t_{j})-\alpha)e^{-\beta_{1}nh}$ to get
$\alpha+(Y(t_{j})-\alpha)e^{-\beta_{1}nh}-Y(t_{j})=(Y(t_{j})-\alpha)(e^{-\beta_{1}nh}-1),$ (2.14)
but this time we need to factorise $e^{-\beta_{1}nh}-1$ into
$e^{-\beta_{1}nh}-1=\Bigg{(}\sum_{i=0}^{n-1}e^{-\beta_{1}ih}\Bigg{)}(e^{-\beta_{1}h}-1),$ (2.15)
which is true because
$\Bigg{(}\sum_{i=0}^{n-1}a^{i}\Bigg{)}(a-1)=\sum_{i=1}^{n}a^{i}-\sum_{i=0}^{n-1}a^{i}=a^{n}-1.$
Substituting Equation (2.15) into Equation (2.14) gives
$\text{LHS}=(Y(t_{j})-\alpha)\Bigg{(}\sum_{i=0}^{n-1}e^{-\beta_{1}ih}\Bigg{)}(e^{-\beta_{1}h}-1).$ (2.16)
Let us return to the right side of Equation (2.13), and look at the argument within the sum. We substitute in the analytic solution to give
$\alpha-Y(t_{j}+ih)=$ $\ \alpha-(\alpha+(Y(t_{j})-\alpha)e^{-\beta_{1}ih})$
$=$ $\ -(Y(t_{j})-\alpha)e^{-\beta_{1}ih}$
$=$ $\ (\alpha-Y(t_{j}))e^{-\beta_{1}ih}$
which we substitute back in to give
$\text{RHS}=f(\hat{\beta}_{1},h)(\alpha-Y(t_{j}))\Bigg{(}\sum_{i=0}^{n-1}e^{-\beta_{1}ih}\Bigg{)}.$ (2.17)
Recombining Equations (2.16) and (2.17), we subtract Equation (2.16) from both sides to give
$0=$ $\ f(\hat{\beta}_{1},h)(\alpha-Y(t_{j}))\Bigg{(}\sum_{i=0}^{n-1}e^{-\beta_{1}ih}\Bigg{)}-(Y(t_{j})-\alpha)\Bigg{(}\sum_{i=0}^{n-1}e^{-\beta_{1}ih}\Bigg{)}(e^{-\beta_{1}h}-1)$
$=$ $\ (\alpha-Y(t_{j}))\Bigg{(}\sum_{i=0}^{n-1}e^{-\beta_{1}ih}\Bigg{)}(f(\hat{\beta}_{1},h)+e^{-\beta_{1}h}-1).$ (2.18)
Equation (2.18) is 0 when
$0=f(\hat{\beta}_{1},h)+e^{-\beta_{1}h}-1=h\hat{\beta}_{1}-\frac{h^{2}}{2}\hat{\beta}_{1}^{2}+\frac{h^{3}}{6}\hat{\beta}_{1}^{3}-\frac{h^{4}}{24}\hat{\beta}_{1}^{4}+e^{-\beta_{1}h}-1,$ (2.19)
which is the same as Equation (2.12) from the 1-step case. As such, $\hat{\beta}_{1}$ solutions will be the same for all $n$. Equation (2.19) does not depend on $t_{j}$, so as long as $Y(t_{j})\neq\alpha$ the solutions are independent of $Y(t_{j})$ as well. ∎
Now that we have a general form for where $\hat{\beta}_{1}$ estimates come from, we can prove non-uniqueness from the polynomial that $\hat{\beta}_{1}$ solves.
Theorem 2.3.
For an affine first order ODE of the form
$\frac{dY}{dt}=\beta_{0}-\beta_{1}Y(t)$
where $\beta_{0}=\alpha\beta_{1}$, within a longitudinal model using $n$ steps of the Runge-Kutta 4th order algorithm, in general there are multiple $\hat{\beta}_{1}$ values that match the analytic solution for sizes over time.
Proof.
From Proposition 2.2 we know that $\hat{\beta}_{1}$ values that produce a numerical projection that matches the analytic solution will solve
$10=h\hat{\beta}_{1}-\frac{h^{2}}{2}\hat{\beta}_{1}^{2}+\frac{h^{3}}{6}\hat{\beta}_{1}^{3}-\frac{h^{4}}{24}\hat{\beta}_{1}^{4}+e^{-\beta_{1}h}-$ 1
which is a polynomial in $\hat{\beta}_{1}$ of degree 4. The solutions can be explicitly calculated using the quartic root formula.
For some parameter values Equation (2.19) may have a combination of real and complex roots, but as there must be at least one solution which approximates with the true (real) parameter $\beta_{1}$, we can exclude the case of all imaginary roots. If the discriminant of the polynomial in $\hat{\beta}_{1}$ is 0 there are some special cases where one real root (of some multiplicity) exists, that can be determined with further calculations when $h$ and $\beta_{1}$ are known. For example, for a given $h$ there will be one $\beta_{1}$ such that there is a single real root of multiplicity four given by
$\hat{\beta}_{1}=-\frac{h^{3}}{6}\bigg{(}\frac{4h^{4}}{24}\bigg{)}^{-1}=-\frac{h^{3}}{6}\frac{24}{4h^{4}}=-\frac{1}{h},$
and $\beta_{1}$ can be determined through calculation using the discriminant when $h$ is chosen (Rees01021922). ∎
One immediate consequence of these results is inaccuracy in the $\hat{\beta}_{1}$ value closest to the true $\beta_{1}$, due to the numerical approximation.
Corollary 2.4.
For an affine first order ODE of the form
$\frac{dY}{dt}=\beta_{0}-\beta_{1}Y(t),\quad\beta_{1}\neq 0$
given true $\beta_{0}$ and $\beta_{1}$, within a longitudinal model using $n$ steps of the Runge-Kutta 4th order algorithm, the $\hat{\beta}_{1}$ closest to the true solution is not equal to it, that is, $\hat{\beta}_{1}\neq\beta_{1}$.
Proof.
We take
$h\hat{\beta}_{1}-\frac{h^{2}}{2}\hat{\beta}_{1}^{2}+\frac{h^{3}}{6}\hat{\beta}_{1}^{3}-\frac{h^{4}}{24}\hat{\beta}_{1}^{4}+e^{-\beta_{1}h}-1$
from Proposition 2.2. Set $\hat{\beta}_{1}=\beta_{1}$ to give
$h\beta_{1}-\frac{h^{2}}{2}\beta_{1}^{2}+\frac{h^{3}}{6}\beta_{1}^{3}-\frac{h^{4}}{24}\beta_{1}^{4}+e^{-\beta_{1}h}-1,$
then we can exploit the series expansion of the exponential function
$e^{x}=\sum_{n=0}^{\infty}\frac{x^{n}}{n!}$
by setting $x=-h\beta_{1}$ to get
$e^{-h\beta_{1}}-\bigg{(}1-h\beta_{1}+\frac{h^{2}\beta_{1}^{2}}{2}-\frac{h^{3}\beta_{1}^{3}}{6}+\frac{h^{4}\beta_{1}^{4}}{24}\bigg{)}=$ $\ \sum_{n=0}^{\infty}\frac{(-h\beta_{1})^{n}}{n!}-\sum_{n=0}^{4}\frac{(-h\beta_{1})^{n}}{n!}$
$=$ $\ \sum_{n=5}^{\infty}\frac{(-h\beta_{1})^{n}}{n!}$
[…]
unless $h$ or $\beta_{1}$ is 0.
As the true value $\hat{\beta_{1}}=\beta_{1}$ does not perfectly match the numeric solution to the analytic one, the estimate closest to $\beta_{1}$ that does solve Equation (2.19) will not be $\hat{\beta_{1}}=\beta_{1}$. ∎
We suspect that Corollary 2.4 implies bias – even for the best of the available estimates of $\beta_{1}$.
We have described an example case where $\beta_{1}=1$, and $\beta_{0}=10$. The two real solutions to Equation (2.19) for those values are $\hat{\beta}_{1}=1.0008$ and $4.9112$, which is shown in Figure 2 and aligns with the behaviour in Figure 1. The small difference from $\beta_{1}=1$ in the better estimate is to be expected given Corollary 2.4.
[…]
To conclude the theory section, the posterior distribution for $\beta_{1}$ and $\beta_{0}$ conditioned on the RK4 step will have modes at the parameter values that produce $\hat{Y}(t_{j})$ that match the analytic solution. The solutions to Equation (2.19) gives posterior modes that cannot be distinguished by estimation based on a longitudinal model using the RK4 algorithm, which is precisely what non-identifiability means in this context.
@@
Equation 2.19 is calculable directly because we have an analytic solution, explicit numerical method, and a straight-forward function. Attempts to do the same with the non-linear Canham function (Canham et al.,, 2004) for example would be impossible as there is no analytic solution for comparison. While the substitution with the analytic solution is not essential to testing a step where a better numerical method can give $Y(t_{j}+h)$, it would be intractable for Canham anyway due to the non-linear DE. In general it may not be possible to do an explicit formula of that form. Simulation is much easier and the user can test different methods and estimation tools as well.
Figure 3: (a) gives the simulated data with measurement error and precision of 0.1 compared to the analytic solution for $\beta_{0}=10$, $\beta_{1}=1$. Outside of the specified situations with independent error for each chain, estimation is based on the same data. (b) is a scatter plot of all posterior estimates, good and bad, for the default priors across three step sizes using RK4. While there are a total of 30 000 points on this plot, there is so little variation in them that they appear at just four points – one near the true value, and for each step size, a second smaller cluster of points stacked around an erroneous value. (c) gives the erroneous differential equations and shows that all estimates converge to the same maximum size. (d) shows the analytic solutions using the erroneous parameter combinations to show that the erroneous values give sizes that converge to the asymptotic maximum much faster than the true solution.
2.2 Aims
We have three questions about the pathological behaviour:
@@ 2.2 Aims
  • Is there a model that provides accurate unimodal estimates? We do simulation studies for the affine and Canham models to test whether the chosen work-arounds – an analytic implementation and RK45 with adaptive step size respectively – have reliable results.
2.3 Methods
To address the questions we simulated data so we know the true values. We produce ‘true’ data from the analytic solution Equation (2.3) with parameters $\beta_{0}=10$, $\beta_{1}=1$ and initial size $Y(0)=1$. The simulated survey structure is 10 size values $Y(t_{j})$ with an interval of 1 unit of time between them, so $t_{j+1}-t_{j}=1$. From the analytic sizes over time we get ‘observations’ with error, rounded to precision 0.1, of the form
@@ 2.3 Methods
Our three questions require a set of empirical tests. For each test we ran 10,000 independent chains on the same dataset. In the case of MCMC this is 10,000 independent chains, while for the optimisation algorithm it is 10,000 runs each of which randomly selects a starting point in the parameter space. The large sample size was chosen because in initial testing we found that the second mode for the smallest RK4 step size was very rare.
@@
From each chain we extract the estimates $\hat{\beta}_{0}$ and $\hat{\beta}_{1}$ as the means of the posterior samples. We fit multivariate normal finite mixture models (McLachlan et al.,, 2019) to the 10,000 estimates for each MCMC model using the mixtools package (Benaglia et al.,, 2009) which provides estimates of cluster means, variances, and mixture probabilities. To initialise the finite mixture algorithm, we classify clusters by taking points on either side of the sample mean for $\hat{\beta}_{0}$, giving
@@
To avoid convergence problems we initialise the finite mixture algorithm with the sample means of the two initial clusters. The fitted mixture of normal distributions is then used to analyse the clustering behaviour, as we get cluster mean estimate vectors, covariance matrices, and cluster occurrence probability estimates. We also use scatter plots to look at the cluster marginal posterior parameter distributions. In the unimodal case we look at the mean of the estimated posterior parameters to check for evidence of large, and a 95% posterior credible interval (Gelman et al.,, 2021) built from taking the central 95% quantile interval of the posterior estimates to get an idea of spread.
To compare the effect of different priors we use the posterior mixtures models and estimates of the posterior density at the posterior modes. We calculate the prior density at each of the posterior modes to see how concentrated the prior is at each point, and for comparison to the posterior probability. If the prior has a large impact we would expect to see posterior probabilities that align with the relative size of the prior density. If the prior is not informative, differences in the prior density will not change the posterior probabilities.
With the estimates in hand we can independently verify the numerical error using the ode function from the deSolve package (Karline Soetaert et al.,, 2010), which has its own implementation of the Runge-Kutta 4th order method (rk4). We plot the $\hat{Y}(t_{j})$ numerical values for the true and estimated parameters with the step size used in the MCMC sampling and calculate the root mean squared error to show that the empirical results reflect the theoretical match between the forward projections. We use plots of the analytic solutions to show that the erroneous parameter combinations actually produce very different behaviour in the absence of numerical error.